The following data were taken from the income statements of Mydorf Company. 2020 2019 Sales revenue $6,370,000 $6,170,000 Beginning inventory 915,000 865,000 Purchases 4,359,750 4,678,000 Ending inventory 1,050,000 915,000 Collapse question part (a1) Compute for each year the inventory turnover. (Round answers to 1 decimal place, e.g. 1.6.) 2020 2019 Inventory turnover times times
Answer #1

Calculation of Inventory Turnover Ratio

COGS / Average inventory = Inventory turnover ratio
2020 (915000+4359750-1050000) = 4224750 / (915000+1050000)/2 = 982500 = 4.3 Times
2109 (865000+4678000-915000) = 4628000 / (865000+915000)/2 = 890000 = 5.2 times